# Olaph: Improving Factuality in Biomedical Long-form Question Answering

### v1

@@ A.2 Experimental Details
We use 8 Nvidia A100 with 80GB memory to train our Olaph framework. Our code is written in PyTorch [67] and HuggingFace [68]. We use Deepspeed stage 3 [69] to implement multi-GPU settings and FlashAttention [70] for efficient training. We use a 5e-7 learning rate with a 0.1 warmup ratio in the initial step and use a 1e-7 learning rate after Step-2 DPO training. We use a 0.01 $\beta$ value to train through DPO learning. For sampling predictions using temperature scaling [32], we generate $k$=6 predictions: one for deterministic prediction ($\tau$ = 0) and five for sampling predictions ($\tau$ = 1.0). To preserve the quality of long-form responses, we set the pre-determined threshold as 200. For inference, we use vllm [71] to speed up our inference time.
A.3 Detail performance of Olaph framework
In Table 4 and 6, we provide the detailed performance used to evaluate three categories: word composition, semantic similarities, and factuality.
@@ A.3 Detail performance of Olaph framework
In detail, word composition consists of R1, R2, and RL scores, each referring to Rouge-1, Rouge-2, and Rouge-L scores [38]. To capture the non-trivial semantic similarities between answer and prediction, we use BLEURT (BL) [41] and BERTScore (BS) [42]. We primarily focus on evaluating factuality automatically using Hallucination (HL) and Comprehensiveness (CP) following previous work [22].
A.4 Biomedical Knowledge Source & Domain-specific Retriever
Data # Documents # Chunks Embedding Size
PubMed 36,533,377 69,743,442 400GB
PMC 1,060,173 46,294,271 160GB
CPG 35,733 606,785 3.5GB
Textbook 18 133,875 0.7GB
We use FActScore [20], which is not used during training, as an additional metric to measure factuality. FActScore measures the support of atomic facts using Wikipedia dumps. However, Wikipedia may not provide sufficient information for discerning biomedical or clinical claims. Therefore, considering the possibility of utilizing a domain-specific knowledge retriever, we follow the construction of biomedical knowledge from documents retrieved by Self-BioRAG [7].
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ A.4 Biomedical Knowledge Source & Domain-specific Retriever
In the fields of medical and clinical domains, researchers and doctors often supplement their knowledge with additional information to address challenging issues effectively. Similarly, for a language model to solve problems, it needs to retrieve relevant documents if necessary. To accomplish this, we utilize the MedCPT [66] retriever off-the-shelf, which is contrastively trained on an unprecedented scale of 255M query-article pairs from PubMed search logs. We compile data from four sources to retrieve relevant documents: PubMed Abstract, PMC Full-text, Clinical Guidelines, and English Medical Textbooks. To ensure computational efficiency, we encode these data offline. The documents are segmented into chunks of 128 words with 32-word overlaps to form evidence, following previous works[72, 73]. Initially, we retrieve the top-20 evidence from each source data, resulting in a total of 80 evidence pieces. Subsequently, we employ the reranking module to obtain the final top-20 evidence relevant to the query. Table 5 presents the overall statistics of the biomedical corpus and the number of indexed documents.
A.5 Why do we use sampling-based predictions?
@@
Dataset: K-QA Golden [22]
Model: BioMistral 7B [6]
@@
Dataset: HealthSearch QA [1]
Model: Self-BioRAG 7B [7]
@@
Dataset: K-QA Silver [22]
Model: Mistral 7B [65]

### v2

@@ A.3 Experimental Details
We use 8 Nvidia A100 with 80GB memory to train our Olaph framework. Our code is written in PyTorch (Paszke et al., 2019) and HuggingFace (Wolf et al., 2019). We use Deepspeed stage 3 (Rajbhandari et al., 2020) to implement multi-GPU settings and FlashAttention (Dao et al., 2022) for efficient training. We use a 5e-7 learning rate with a 0.1 warmup ratio in the initial step and use a 1e-7 learning rate after Step-2 DPO training. We use a 0.01 $\beta$ value to train through DPO learning. For sampling predictions using temperature sampling (Guo et al., 2017), we generate $k$=6 predictions: one for deterministic prediction ($\tau$ = 0) and five for sampling predictions ($\tau$ = 1.0). To preserve the quality of long-form responses, we set the pre-determined threshold as 200. For inference, we use vllm (Kwon et al., 2023) to speed up our inference time.
@@
A.4 Detail performance of Olaph framework
In Table 8 and 9, we provide the detailed performance used to evaluate three categories: word composition, semantic similarities, and factuality. In detail, word composition consists of R1, R2, and RL scores, each referring to Rouge-1, Rouge-2, and Rouge-L scores (Lin, 2004). To capture the non-trivial semantic similarities between answer and prediction, we use BLEURT (BL) (Sellam et al., 2020) and BERTScore (BS) (Zhang et al., 2019). We primarily focus on evaluating factuality automatically using Hallucination (HL) and Comprehensiveness (CP) following previous work (Manes et al., 2024).
A.5 Biomedical Knowledge Source & Domain-specific Retriever
Data # Documents # Chunks Embedding Size
PubMed 36,533,377 69,743,442 400GB
PMC 1,060,173 46,294,271 160GB
CPG 35,733 606,785 3.5GB
Textbook 18 133,875 0.7GB
We use FActScore (Min et al., 2023), which is not used during training, as an additional metric to measure factuality. FActScore measures the support of atomic facts using Wikipedia dumps. However, Wikipedia may not provide sufficient information for discerning biomedical or clinical claims. Therefore, considering the possibility of utilizing a domain-specific knowledge retriever, we follow the construction of biomedical knowledge from documents retrieved by Self-BioRAG (Jeong et al., 2024).
@@ A.5 Biomedical Knowledge Source & Domain-specific Retriever
In the fields of medical and clinical domains, researchers and doctors often supplement their knowledge with additional information to address challenging issues effectively. Similarly, for a language model to solve problems, it needs to retrieve relevant documents if necessary. To accomplish this, we utilize the MedCPT (Jin et al., 2023) retriever off-the-shelf, which is contrastively trained on an unprecedented scale of 255M query-article pairs from PubMed search logs. We compile data from four sources to retrieve relevant documents: PubMed Abstract, PMC Full-text, Clinical Guidelines, and English Medical Textbooks. To ensure computational efficiency, we encode these data offline. The documents are segmented into chunks of 128 words with 32-word overlaps to form evidence, following previous works(Wang et al., 2019; Karpukhin et al., 2020). Initially, we retrieve the top-20 evidence from each source data, resulting in a total of 80 evidence pieces. Subsequently, we employ the reranking module to obtain the final top-20 evidence relevant to the query. Table 10 presents the overall statistics of the biomedical corpus and the number of indexed documents.
A.6 Details of Entailment Model
We employed hallucination and comprehensiveness metrics to evaluate factuality in an automated and cost-effective manner. This involved assessing the degree of entailment, specifically how much two statements are included in the actual model’s response. Further details about this model will be provided in the appendix of the revised paper. In Table 11, we use a model fine-tuned on BioBERT (Lee et al., 2020) with three NLI datasets, MultiNLI (Williams et al., 2018), SNLI (Bowman et al., 2015), and MedNLI (Romanov & Shivade, 2018). These datasets are designed to determine the inference relationship between two texts. The table below presents the performance (i.e., accuracy) of the test sets for the two datasets used to train the model, as well as the performance on MedNLI, which is used in the medical domain. While easy-to-access models like GPT-3.5 or GPT-4 could be used for entailment, we aimed to utilize models that are widely-used to many medical researchers without incurring significant API costs.
Table 11: Performance of Entailment Model in NLI datasets.
Model MultiNLI-Matched MultiNLI-Mismatched SNLI MedNLI
GPT-3.5-turbo 69.4 69.3 65.7 59.2
BioBERT-NLI 89.3 88.8 89.2 85.5
A.7 Why do we use sampling-based predictions?
@@
Dataset: K-QA Golden (Manes et al., 2024)
Model: BioMistral 7B (Labrak et al., 2024)
@@
Dataset: HealthSearch QA (Singhal et al., 2023)
Model: Self-BioRAG 7B (Jeong et al., 2024)
@@
Dataset: K-QA Silver (Manes et al., 2024)
Model: Mistral 7B (Jiang et al., 2023)